# Revealing POMDPs: Qualitative and Quantitative Analysis for Parity Objectives

### v1

@@ 1 Introduction
Partially observable Markov decision processes (POMDPs) model sequential decision-making under uncertainty (Bertsekas 1976; Papadimitriou and Tsitsiklis 1987; Kaelbling, Littman, and Cassandra 1998). At each step, the environment is in some hidden state. A controller interacts with it by choosing actions. The chosen action and the current state determine a probability distribution over the subsequent state. The controller cannot observe the state directly, but observes a signal that discloses only partial information of the state. POMDPs generalize classic models: Markov decision processes (MDPs), in which the state is fully observed (Puterman 2014), and blind MDPs, in which no state information is observed and are equivalent to Probabilistic Finite Automata (Rabin 1963; Paz 1971).
@@
The controller aims to maximize an objective function, which formally captures the desired behaviors of the model. Two main classes of objectives are typically considered: logical objectives, e.g., reachability and LTL (linear-temporal logic), and quantitative objectives, e.g., discounted sum and limit-average; see (Puterman 2014; Baier and Katoen 2008; Filar and Vrieze 1997) for details. This work focuses on logical objectives.
The most basic example of logical objectives is reachability, i.e., given a set of target states, the objective requires that some target state is visited at least once. A more general class of logical objectives is parity objectives, which assign to each state a non-negative integer, called priority. The objective is satisfied if the smallest priority appearing infinitely often is even. Parity objectives express all commonly used temporal objectives such as liveness, and are a canonical form to express all $\omega$-regular objectives (Thomas 1997) and LTL objectives (Baier and Katoen 2008). Hence, the study of POMDPs with parity objectives is a fundamental theoretical problem.
@@
POMDPs have been applied across diverse fields, including computational biology (Durbin et al. 1998) and reinforcement learning (Kaelbling, Littman, and Moore 1996). In particular, POMDPs with logical objectives have proven useful in many application areas such as probabilistic planning (Bonet and Geffner 2009); randomized embedded scheduler (de Alfaro et al. 2005); randomized distributed algorithms (Pogosyants, Segala, and Lynch 2000); probabilistic specification languages (Baier, Größer, and Bertrand 2012); and robot planning (Kress-Gazit, Fainekos, and Pappas 2009; Chatterjee et al. 2015).
@@
Most results for POMDPs and the computational analysis are negative (undecidability results). Indeed, the limit-sure analysis for reachability objectives is undecidable (Gimbert and Oualhadj 2010; Chatterjee and Henzinger 2010), which extends to general parity objectives. The almost-sure analysis for reachability objectives is EXPTIME-complete (Baier, Bertrand, and Größer 2008; Chatterjee, Doyen, and Henzinger 2010), but is undecidable for parity objectives even for two priorities (namely coBüchi objectives) (Baier, Bertrand, and Größer 2008; Chatterjee et al. 2010). The quantitative analysis for reachability objectives is undecidable (Madani, Hanks, and Condon 2003), which extends to general parity objectives. Given this wide range of results about non-existence of algorithms a natural direction to explore is the existence of subclasses for which the computational problems are decidable.
@@
We study revealing POMDPs, a special subclass of POMDPs where each visited state is announced to the controller with a positive probability. Consequently, the controller’s uncertainty about the state, i.e. the probability distribution over states (the belief), occasionally collapses into a Dirac distribution on the announced state. Revealing POMDPs have been previously studied in several works, including (Chen and Liew 2023; Belly et al. 2025; Avrachenkov, Dhiman, and Kavitha 2025), which present motivation and applications for this model.
@@
A key result for revealing POMDPs shows that the almost-sure analysis for parity objectives is EXPTIME-complete (Belly et al. 2025). However, the limit-sure and quantitative analysis problems for reachability and parity objectives remained open for revealing POMDPs.
@@
A closely related work established that almost-sure analysis for parity objectives in revealing POMDPs is EXPTIME-complete (Belly et al. 2025). Additionally, (Chatterjee, Chmelík, and Tracol 2016) previously showed that almost-sure analysis for parity objectives in general POMDPs under finite-memory policies is EXPTIME-complete. Therefore, the EXPTIME-completeness result naturally follows once finite memory policies are proven sufficient for almost-sure analysis. However, limit-sure analysis and quantitative analysis for POMDPs remain undecidable in general, even under finite memory policies (Chatterjee, Saona, and Ziliotto 2021). This highlights a sharp contrast with the almost-sure analysis and gives rise to new technical challenges.
@@
The study of subclasses of POMDPs with tractable algorithms is a broad topic with many directions. For example, subclasses of POMDPs for qualitative analysis have been explored in (Fijalkow et al. 2015; Chatterjee and Tracol 2012); and for quantitative analysis various subclasses have also been studied such as with ergodicity condition (Chatterjee et al. 2024); multiple environments only (Van Der Vegt, Jansen, and Junges 2023; Chatterjee et al. 2025); or in online learning (Liu et al. 2022; Chen et al. 2023). Our work focuses on such a subclass, namely revealing POMDPs, which has been studied in the literature.
Avrachenkov, Dhiman, and Kavitha (2025) studied strongly-connected revealing POMDPs and restricting attention to the set of belief-stationary policies. They proved that there is an optimal policy within this set, whose induced belief dynamics are contracting and reach a positive recurrent class of beliefs in finite time. They also provided an infinite-dimensional linear program that characterizes the optimal belief-stationary policy. They considered an application of strongly-connected revealing blind MDPs. (Chen and Liew 2023) introduced the model of revealing blind MDPs. They consider the discounted objective and present algorithms based on finite-horizon approximation to compute the discounted value. Our work addresses the open computational analysis questions for revealing POMDPs.
@@ A POMDP is a tuple $P=(\mathcal{S},\mathcal{A},\mathcal{Z},\delta,b_{0})$, where:
Markov Decision Processes (MDPs) are POMDPs in which the signal corresponds to the state (Puterman 2014). Formally, an MDP is denoted by $M=(\mathcal{S},\mathcal{A},\delta,b_{0})$ with transition function $\delta\colon\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$.
@@
Revealing POMDPs coincide with the class of strongly revealing defined in (Belly et al. 2025). Indeed, if the controller observes a signal $s\in\mathcal{Z}$, then their next belief is $\mathds{1}[s]$, which corresponds to a revelation of the state.
@@
  • In general POMDPs, it is enough to consider transition functions that pair each state with only one signal, known as deterministic observation, see (Chatterjee, Chmelík, and Tracol 2016, Remark 4). For revealing POMDPs, deterministic observations do not capture the general case because they correspond to MDPs.
@@
For a finite prefix of a play, an element in $(\mathcal{S}\times\mathcal{A})^{*}$, its cone is the set of plays with it as their prefix. Given a policy $\sigma$ and an initial belief $b_{0}$, the unique probability measure over Borel sets of infinite plays obtained given $\sigma$ is denoted by $\mathbb{P}^{\sigma}_{b_{0}}(\cdot)$, which is defined by Carathéodory’s extension theorem by extending the natural definition over cones of plays (Billingsley 2012).
@@
An objective in a POMDP is a Borel set of plays $\Phi\subseteq\Omega$ in the Cantor topology on $\Omega$ (Kechris 1995). We consider objectives in the first $2\nicefrac{{1}}{{2}}$ levels of the Borel hierarchy, including the parity objective which expresses all $\omega$-regular objectives (Thomas 1997). Denote the state at time $t$ by $s_{t}$ and set of states that occur infinitely often in a play $\rho$ by $\mathcal{I}(\rho)\coloneqq\{s\in\mathcal{S}:\forall t\geq 0\;\exists\widetilde{t}\geq t\quad s=s_{\widetilde{t}}\in\rho\}$. We consider the following objectives.
  • Reachability: Given a set $\mathcal{X}\subseteq\mathcal{S}$ of target states, the reachability objective requires that a target state is visited at least once. Formally, the reachability objective is $\operatorname{Reach}(\mathcal{X})\coloneqq\{\rho\in\Omega:\exists t\geq 0\quad s_{t}\in\mathcal{X}\}$.
  • Parity: Given a $d\geq 0$ and a function $\operatorname{pri}\colon\mathcal{S}\to\{0,1,\ldots,d\}$ of priorities, the parity objective requires that the smallest priority that appears infinitely often is even. Formally, $\operatorname{Parity}\coloneqq\left\{\rho\in\Omega:\min\{\operatorname{pri}(s):s\in\mathcal{I}(\rho)\}\text{ is even}\right\}$.
@@ 3.2 Overview of Results
The following results follow from (Belly et al. 2025, Theorem 3) and Theorem 4.
@@ 4 Belief-Reachability Objectives
Similar to (Andersson and Miltersen 2009, Lemma 5) that considers finite duration games, it is easy to see that, if a policy is stopping and uses only reliable actions, then it is optimal.
@@
$\delta_{\min}\coloneqq\min\Bigl\{\delta(s,a)(s^{\prime},z)\;:\;\forall s,\,s^{\prime}\in\mathcal{S},\;a\in\mathcal{A},\;z\in\mathcal{Z}$
$\delta(s,a)(s^{\prime},z)>0\Bigr\}.$
@@
Finite horizon objectives have been studied by a long time. A naive approach to the quantitative analysis for belief-reachability objectives would take the time bound $T$ in Lemma 10 and solve a (fully observable) MDP with $|\mathcal{A}\times\mathcal{S}|^{T}$ states. This approach implies a 2EXPTIME complexity upper bound. A fundamental result states that, for horizons that are polynomially large with respect to the input, computing the finite horizon value of POMDPs is PSPACE-complete (Papadimitriou and Tsitsiklis 1987, Theorem 6, page 448). The technique, instead of listing explicitly all exponentially many histories, compactly represents them using nondeterminism and space proportional to the horizon. The conclusion follows from PSPACE being closed under nondeterminism by Savitch’s theorem. For exponentially large horizons, this technique leads to an EXPSPACE upper bound. Instead, we prove an EXPTIME upper bound.
Point-based algorithms were introduced for POMDPs by (Pineau, Gordon, and Thrun 2003) as an alternative to approximating the value of POMDPs by considering a fixed subset of beliefs and projected belief updates. For a survey on point-based algorithms see (Shani, Pineau, and Kaplow 2013; Walraven and Spaan 2019), which focuses on finite horizon objectives but does not provide the complexity upper bound we require.
@@ 5.1 Reduction to Belief-Reachability Objectives
In the sequel, we show that, for revealing POMDPs, the parity objective value coincides with the belief-reachability to the set of Dirac on the almost‐sure winning parity states. This proof uses the standard notion of end components in MDPs, introduced in (Alfaro 1998); see also (Baier and Katoen 2008, Section 10.6.3).
@@
Lemma 12 ((Alfaro 1998, Theorems 3.1 and 3.2)).
@@
Consider the example presented in (Chatterjee, Saona, and Ziliotto 2021, Example 4.3) of a POMDP with only one possible signal in Figure 1, commonly known as a blind MDP. It has four states $\mathcal{S}=\{s_{0},s_{1},\bot,\top\}$ and two actions: wait ($w$) and commit ($c$), i.e., $\mathcal{A}=\{w,c\}$. The priority function is defined as
@@ 5.2 Proofs of Main Results
By (Belly et al. 2025, Theorem 3), computing almost-sure winning parity states in the revealing POMDP is EXPTIME. Therefore, this result is a direct implication of Lemma 15 and Theorem 1. ∎
@@ Appendix B Proofs of Section 5
By (Belly et al. 2025, Theorem 3), computing almost-sure winning parity states in the revealing POMDP is in EXPTIME. By Theorem 1, approximating the belief-reachability value to this set of states is in EXPTIME. By Lemma 15, this value coincides with the parity value and therefore we computed an approximation of the parity value in EXPTIME. ∎
@@
Consider revealing POMDPs with parity objectives. By Theorem 4, optimal policies exist. Therefore, if a revealing POMDP with parity objectives is limit-sure winning, then it is almost-sure winning. The reverse direction follows by definition. Lastly, by (Belly et al. 2025, Theorem 3), the existence of an almost-sure policy is in EXPTIME-complete. Therefore, limit-sure analysis is also EXPTIME-complete because the two problems coincide. ∎

### v2

@@ 1 Introduction
Partially observable Markov decision processes (POMDPs) model sequential decision-making under uncertainty Bertsekas (1976); Papadimitriou and Tsitsiklis (1987); Kaelbling, Littman, and Cassandra (1998). At each step, the environment is in some hidden state. A controller interacts with it by choosing actions. The chosen action and the current state determine a probability distribution over the subsequent state. The controller cannot observe the state directly, but observes a signal that discloses only partial information of the state. POMDPs generalize classic models: Markov decision processes (MDPs), in which the state is fully observed Puterman (2014), and blind MDPs, in which no state information is observed and are equivalent to Probabilistic Finite Automata Rabin (1963); Paz (1971).
@@
The controller aims to maximize an objective function, which formally captures the desired behaviors of the model. Two main classes of objectives are typically considered: logical objectives, e.g., reachability and LTL (linear-temporal logic), and quantitative objectives, e.g., discounted sum and limit-average; see Puterman (2014); Baier and Katoen (2008); Filar and Vrieze (1997) for details. This work focuses on logical objectives.
The most basic example of logical objectives is reachability, i.e., given a set of target states, the objective requires that some target state is visited at least once. A more general class of logical objectives is parity objectives, which assign to each state a non-negative integer, called priority. The objective is satisfied if the smallest priority appearing infinitely often is even. Parity objectives express all commonly used temporal objectives such as liveness, and are a canonical form to express all $\omega$-regular objectives Thomas (1997) and LTL objectives Baier and Katoen (2008). Hence, the study of POMDPs with parity objectives is a fundamental theoretical problem.
@@
POMDPs have been applied across diverse fields, including computational biology Durbin et al. (1998) and reinforcement learning Kaelbling, Littman, and Moore (1996). In particular, POMDPs with logical objectives have proven useful in many application areas such as probabilistic planning Bonet and Geffner (2009); randomized embedded scheduler de Alfaro et al. (2005); randomized distributed algorithms Pogosyants, Segala, and Lynch (2000); probabilistic specification languages Baier, Größer, and Bertrand (2012); and robot planning Kress-Gazit, Fainekos, and Pappas (2009); Chatterjee et al. (2015).
@@
Most results for POMDPs and the computational analysis are negative (undecidability results). Indeed, the limit-sure analysis for reachability objectives is undecidable Gimbert and Oualhadj (2010); Chatterjee and Henzinger (2010), which extends to general parity objectives. The almost-sure analysis for reachability objectives is EXPTIME-complete Baier, Bertrand, and Größer (2008); Chatterjee, Doyen, and Henzinger (2010), but is undecidable for parity objectives even for two priorities (namely coBüchi objectives) Baier, Bertrand, and Größer (2008); Chatterjee et al. (2010). The quantitative analysis for reachability objectives is undecidable Madani, Hanks, and Condon (2003), which extends to general parity objectives. Given this wide range of results about non-existence of algorithms a natural direction to explore is the existence of subclasses for which the computational problems are decidable.
@@
We study revealing POMDPs, a special subclass of POMDPs where each visited state is announced to the controller with a positive probability. Consequently, the controller’s uncertainty about the state, i.e. the probability distribution over states (the belief), occasionally collapses into a Dirac distribution on the announced state. Revealing POMDPs have been previously studied in several works, including Chen and Liew (2023); Belly et al. (2025); Avrachenkov, Dhiman, and Kavitha (2025), which present motivation and applications for this model.
@@
A key result for revealing POMDPs shows that the almost-sure analysis for parity objectives is EXPTIME-complete Belly et al. (2025). However, the limit-sure and quantitative analysis problems for reachability and parity objectives remained open for revealing POMDPs.
@@
A closely related work established that almost-sure analysis for parity objectives in revealing POMDPs is EXPTIME-complete Belly et al. (2025). Additionally, Chatterjee, Chmelík, and Tracol (2016) previously showed that almost-sure analysis for parity objectives in general POMDPs under finite-memory policies is EXPTIME-complete. Therefore, the EXPTIME-completeness result naturally follows once finite memory policies are proven sufficient for almost-sure analysis. However, limit-sure analysis and quantitative analysis for POMDPs remain undecidable in general, even under finite memory policies Chatterjee, Saona, and Ziliotto (2021). This highlights a sharp contrast with the almost-sure analysis and gives rise to new technical challenges.
@@
The study of subclasses of POMDPs with tractable algorithms is a broad topic with many directions. For example, subclasses of POMDPs for qualitative analysis have been explored in Fijalkow et al. (2015); Chatterjee and Tracol (2012); and for quantitative analysis various subclasses have also been studied such as with ergodicity condition Chatterjee et al. (2024); multiple environments only Van Der Vegt, Jansen, and Junges (2023); Chatterjee et al. (2025); or in online learning Liu et al. (2022); Chen et al. (2023). Our work focuses on such a subclass, namely revealing POMDPs, which has been studied in the literature.
Avrachenkov, Dhiman, and Kavitha (2025) studied strongly-connected revealing POMDPs and restricting attention to the set of belief-stationary policies. They proved that there is an optimal policy within this set, whose induced belief dynamics are contracting and reach a positive recurrent class of beliefs in finite time. They also provided an infinite-dimensional linear program that characterizes the optimal belief-stationary policy. They considered an application of strongly-connected revealing blind MDPs. Chen and Liew (2023) introduced the model of revealing blind MDPs. They consider the discounted objective and present algorithms based on finite-horizon approximation to compute the discounted value. Our work addresses the open computational analysis questions for revealing POMDPs.
@@ A POMDP is a tuple $P=(\mathcal{S},\mathcal{A},\mathcal{Z},\delta,b_{0})$, where:
Markov Decision Processes (MDPs) are POMDPs in which the signal corresponds to the state Puterman (2014). Formally, an MDP is denoted by $M=(\mathcal{S},\mathcal{A},\delta,b_{0})$ with transition function $\delta\colon\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$.
@@
Revealing POMDPs coincide with the class of strongly revealing defined in Belly et al. (2025). Indeed, if the controller observes a signal $s\in\mathcal{Z}$, then their next belief is $\mathds{1}[s]$, which corresponds to a revelation of the state.
@@
  • In general POMDPs, it is enough to consider transition functions that pair each state with only one signal, known as deterministic observation, see (Chatterjee, Chmelík, and Tracol, 2016, Remark 4). For revealing POMDPs, deterministic observations do not capture the general case because they correspond to MDPs.
@@
For a finite prefix of a play, an element in $(\mathcal{S}\times\mathcal{A})^{*}$, its cone is the set of plays with it as their prefix. Given a policy $\sigma$ and an initial belief $b_{0}$, the unique probability measure over Borel sets of infinite plays obtained given $\sigma$ is denoted by $\mathbb{P}^{\sigma}_{b_{0}}(\cdot)$, which is defined by Carathéodory’s extension theorem by extending the natural definition over cones of plays Billingsley (2012).
@@
An objective in a POMDP is a Borel set of plays $\Phi\subseteq\Omega$ in the Cantor topology on $\Omega$ Kechris (1995). We consider objectives in the first $2\nicefrac{{1}}{{2}}$ levels of the Borel hierarchy, including the parity objective which expresses all $\omega$-regular objectives Thomas (1997). Denote the state at time $t$ by $s_{t}$ and set of states that occur infinitely often in a play $\rho$ by $\mathcal{I}(\rho)\coloneqq\{s\in\mathcal{S}:\forall t\geq 0\;\exists\widetilde{t}\geq t\quad s=s_{\widetilde{t}}\in\rho\}$. We consider the following objectives.
  • Reachability: Given a set $\mathcal{X}\subseteq\mathcal{S}$ of target states, the reachability objective requires that a target state is visited at least once. Formally, the reachability objective is
    $\operatorname{Reach}(\mathcal{X})\coloneqq\{\rho\in\Omega:\exists t\geq 0\quad s_{t}\in\mathcal{X}\}.$
  • Parity: Given a $d\geq 0$ and a function $\operatorname{pri}\colon\mathcal{S}\to\{0,1,\ldots,d\}$ of priorities, the parity objective requires that the smallest priority that appears infinitely often is even. Formally,
    $\operatorname{Parity}\coloneqq\left\{\rho\in\Omega:\min\{\operatorname{pri}(s):s\in\mathcal{I}(\rho)\}\text{ is even}\right\}.$
@@ 3.2 Overview of Results
The following results follow from (Belly et al., 2025, Theorem 3) and Theorem 4.
@@ 4 Belief-Reachability Objectives
Similar to (Andersson and Miltersen, 2009, Lemma 5) that considers finite duration games, it is easy to see that, if a policy is stopping and uses only reliable actions, then it is optimal.
@@
$\delta_{\min}\coloneqq\min\Bigl\{\delta(s,a)(s^{\prime},z)\;:\;\forall s,\,s^{\prime}\in\mathcal{S},\;a\in\mathcal{A},\;z\in\mathcal{Z}\quad\delta(s,a)(s^{\prime},z)>0\Bigr\}.$
@@
Finite horizon objectives have been studied by a long time. A naive approach to the quantitative analysis for belief-reachability objectives would take the time bound $T$ in Lemma 10 and solve a (fully observable) MDP with $|\mathcal{A}\times\mathcal{S}|^{T}$ states. This approach implies a 2EXPTIME complexity upper bound. A fundamental result states that, for horizons that are polynomially large with respect to the input, computing the finite horizon value of POMDPs is PSPACE-complete (Papadimitriou and Tsitsiklis, 1987, Theorem 6, page 448). The technique, instead of listing explicitly all exponentially many histories, compactly represents them using nondeterminism and space proportional to the horizon. The conclusion follows from PSPACE being closed under nondeterminism by Savitch’s theorem. For exponentially large horizons, this technique leads to an EXPSPACE upper bound. Instead, we prove an EXPTIME upper bound.
Point-based algorithms were introduced for POMDPs by Pineau, Gordon, and Thrun (2003) as an alternative to approximating the value of POMDPs by considering a fixed subset of beliefs and projected belief updates. For a survey on point-based algorithms see Shani, Pineau, and Kaplow (2013); Walraven and Spaan (2019), which focuses on finite horizon objectives but does not provide the complexity upper bound we require.
@@ 5.1 Reduction to Belief-Reachability Objectives
In the sequel, we show that, for revealing POMDPs, the parity objective value coincides with the belief-reachability to the set of Dirac on the almost‐sure winning parity states. This proof uses the standard notion of end components in MDPs, introduced in Alfaro (1998); see also (Baier and Katoen, 2008, Section 10.6.3).
@@
Lemma 12 ((Alfaro, 1998, Theorems 3.1 and 3.2)).
@@
Consider the example presented in (Chatterjee, Saona, and Ziliotto, 2021, Example 4.3) of a POMDP with only one possible signal in Figure 1, commonly known as a blind MDP. It has four states $\mathcal{S}=\{s_{0},s_{1},\bot,\top\}$ and two actions: wait ($w$) and commit ($c$), i.e., $\mathcal{A}=\{w,c\}$. The priority function is defined as
@@ 5.2 Proofs of Main Results
By (Belly et al., 2025, Theorem 3), computing almost-sure winning parity states in the revealing POMDP is EXPTIME. Therefore, this result is a direct implication of Lemma 15 and Theorem 1. ∎
@@ Appendix B Proofs of Section 5
By [7, Theorem 3], computing almost-sure winning parity states in the revealing POMDP is in EXPTIME. By Theorem 1, approximating the belief-reachability value to this set of states is in EXPTIME. By Lemma 15, this value coincides with the parity value and therefore we computed an approximation of the parity value in EXPTIME. ∎
@@
Consider revealing POMDPs with parity objectives. By Theorem 4, optimal policies exist. Therefore, if a revealing POMDP with parity objectives is limit-sure winning, then it is almost-sure winning. The reverse direction follows by definition. Lastly, by [7, Theorem 3], the existence of an almost-sure policy is in EXPTIME-complete. Therefore, limit-sure analysis is also EXPTIME-complete because the two problems coincide. ∎